OpenStudy (anonymous):

A 40.6 g sample of iron metal is heated and put into 100 g of water at 19.7 degrees Celsius in a calorimeter. If the final Temperature of the iron sample and the water is 24.3 degrees Celsius, what was the temperature of the iron sample when it was placed in the water?(I'll give a medal)

OpenStudy (zale101):

\(\LARGE -q_{~lost ~by ~iron} = q_{~gained~by~ water} \) \(~~~~~~~~~\large \color{purple}{-m*C*\Delta T =m*C*\Delta T} \)

OpenStudy (zale101):

m represents the mass, C represents the specific heat capacity, delta T represents the temperature difference/change Keep in mind that \(\large \Delta T=T_{final}−T_{initial}\) You are looking for T initial of iron and the rest is giving to you. Plug everything and solve T initial of Fe algebraically. −m∗C∗ΔT=m∗C∗ΔT \(\large −m_{Fe}∗C_{Fe}∗(T_f-\color{red}{T_i})=m_{H_2O}∗C_{H_2O}∗(T_f-T_i)\)

OpenStudy (zale101):

Mass is always in grams

OpenStudy (zale101):

and again, it 's already been given to you for both iron and water :)

OpenStudy (zale101):

Specific heat capacity should be given to you specific heat C. for iron = 0.44 J/g°C specific heat C. for water = 4.18 J/g°C

OpenStudy (zale101):

Temperatures are already given to you, expect for T initial for iron and that's what you are solving for.